A rectangular pool is 4m wide and 8m long. It is surrounded by a concrete walkway of uniform width. The cost of the walkway is dependent upon its area.
Write a simplified expression representing the total area of the walkway in terms of its width.

Answer by mananth(16946)

pool 4 by 8
area = 32 m^2
..
let width be x
length with walkway = 8+2x
width with walkaway = 4+2x
Total area = (8+2x)(4+2x)
Total area - area of pool = area of walkaway
(8+2x)(4+2x)-32= walkaway area
32+16x+8x+4x^2-32
4x^2+24x = Area of walkaway.
...
m.ananth@hotmail.ca